Calculate the location of Ei in silicon at liquid nitrogen temperature (77 K), at room



temperature (300 K), and at 100 °C (let mp =1.0 m0 and mn= 0.19 m0). Is it reasonable to assume



that Ei is in the center of the forbidden gap?




1
Expert's answer
2022-02-09T16:15:19-0500

"E=\frac{kT}2\ln(\frac{m_p}{m_n}),"

a)

"E_1=88.2\cdot 10^{-23}~J,"

b)

"E_2=343.8\cdot 10^{-23}~J,"

c)

"E_3=427.4\cdot 10^{-23}~J."